OpenStudy (anonymous):

Using Gauss approach a) 1+2+3+4+...+98 I think the answer is 4851 b)1+3+5+7+...+1003 I got two answers for 503506 I wanted to make sure I did this correctly I did 98*99/2 for the first one and 1003*1004/2 for the second on is right?

OpenStudy (luigi0210):

Welcome to Openstudy :)

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

I am not sure if I did that correctly and the first answer I got for b was 1006009 but the second following a previous question I got 503506 for b. Is my second answer correct and if not can you show me how to get the correct answer?

OpenStudy (anonymous):

First is correct .For second see below a=1,d=3-1=2,l=1003 i=a+(n-1) d 1003=1+(n-1)*2 \[n-1=\frac{ 1002 }{ 2 }=501,n=501+1=502\] formula is \[\frac{ n }{ 2 }\left( a+l \right)\] where a is first term and l is last term,n=number of terms.

OpenStudy (anonymous):

I think I got it I went back over and my answer was 252004 because 502th stage is 1003

OpenStudy (anonymous):

\[1+2+3+4+...+98=\frac{98\times 99}{2}\] for the first one

OpenStudy (anonymous):

4851 for the first one

OpenStudy (anonymous):

second one is \(n^2\) were \(n\) is the number of odd numbers you have you can find \(n\) by solving \[2n-1=1003\]

OpenStudy (anonymous):

252004 for b